. Consider the function f(x) = 2x^ 3 + 9x^ 2 − 24x + 1. (1) (5 points) Find all the critical points of f(x).

(2) (5 points) Find the intervals on which f(x) is increasing and decreasing, and find the x-values of any relative minima/maxima.
Answer #1

Dm 1. Given foxsa 23+98-24x+1 :) Fox Critical points: +(22=0 (222792²- 24x+10=0 (3) = me 6x2 + 180-24617=0 6(x²+3x-4)=0 (x +